Can anyone help with this problem
Which of the ordered pairs (-4, 2), (6,)), (-9, -3), (-14, 4) are solutions for the equation x+5y=6 and show work?

Answer by Earlsdon(6294)

I'll show you how to do one then you can finish the rest yourself:
Let's try the first ordered pair (-4, 2).
Remember that an ordered pair gives you the x- and y-coordinates of a point and is always written thus: (x, y)
So the ordered pair (-4, 2) means that the given point has an x-coordinate of -4 and a y-coordinate of 2.
To see if this point is a solution for (satisfies) the equation: x+5y = 6, you simply substitute the x- and y-values from the ordered into the equation.
If the two sides are equal, then the ordered pair is a solution for that equation. If not, then it isn't!
Substitute x = -4 and y = 2 into x + 5y = 6
-4 + 5(2) = 6 ?
-4 + 10 = 6 ?
6 = 6 BINGO! The ordered pair (-4, 2) is a solution!